3u^2 - 5 = 14u
can someone please solve ASAP

Answer by rothauserc(4718)

we are given 3u^2 - 5 = 14u
subtract 14u from both sides of the = and we get
3u^2 - 14u - 5 = 0
this can be factored into the product of
(3u + 1) * (u - 5) = 0
and u = 5 or u = -1/3